At present, Gilbert's age is 30% of his father's age. Thirty years from now, Gilbert's age will be 60% of his father's age. How old is Gilbert now?
Found 2 solutions by ankor@dixie-net.com, aarthi089@gmail.com:
Answer by ankor@dixie-net.com(22740)

At present, Gilbert's age is 30% of his father's age.
g = .3f (using the decimal equiv of percent)
:
Thirty years from now, Gilbert's age will be 60% of his father's age.
g + 30 = .6(f+30)
g + 30 = .6f + 18
g = .6f = 18 - 30
g = .6f - 12
Replace g with .3f
.3f = .6f - 12
12 = .6f - .3f
12 = .3f
f = 12/.3
f = 40 yrs is Dad's age
How old is Gilbert now?
g = .3(40)
g = 12 yrs
:
:
Check this
12 + 30 = .6(40+30)
42 = .6(70)

Answer by aarthi089@gmail.com(1)

cant undrstnd step 3..
pplease explain in detail,,,